# Avatar Concept Slider: Manipulate Concepts In Your Human Avatar With Fine-grained Control

### v2

@@ 4.1 Precise Slider-like Concept Editing
$\mathcal{L}_{sliding}(\Delta\boldsymbol{\phi})=\mathbb{E}_{\alpha\in[-1,1]}\Big{[}\big{|}\big{|}\frac{\mathbf{f}(\alpha,\Delta\boldsymbol{\phi},c_{nu})\cdot\mathbf{b}_{c}}{\mathbf{b}_{c}\cdot\mathbf{b}_{c}}\mathbf{b}_{c}^{T}-\big{(}\frac{1+\alpha}{2}\boldsymbol{\mu}_{p}+\frac{1-\alpha}{2}\boldsymbol{\mu}_{n}\big{)}\big{|}\big{|}\Big{]},$ (3)
@@
where $\mathbf{f}(\cdot)$ represents the generation of latent features, $c_{nu}$ is a neutral text prompt template, |||||$\cdot$||| | ⋅ | | represents L2 distance, $\boldsymbol{\mu}_{p},\boldsymbol{\mu}_{n}\in\mathbb{R}^{D}$ are computed mean vectors for the positive and negative concept sides (see Supp for more details), and various $\alpha$ values are sampled within $[-1,1]$ to represent intermediate concepts. Intuitively, $\frac{\mathbf{f}(\alpha,\Delta\boldsymbol{\phi},c_{nu})\cdot\mathbf{b}_{c}^{T}}{\mathbf{b}_{c}\cdot\mathbf{b}_{c}}\mathbf{b}_{c}^{T}$ is the projection of $\mathbf{f}(\alpha,\Delta\boldsymbol{\phi},c_{nu})$ onto the concept axis $\mathbf{b}_{c}$, and this loss tries to match the projection’s position with the desired location along the concept axis, which is the interpolation between $\boldsymbol{\mu}_{n}$ and $\boldsymbol{\mu}_{p}$ as: $\big{(}\frac{1+\alpha}{2}\boldsymbol{\mu}_{p}+\frac{1-\alpha}{2}\boldsymbol{\mu}_{n}\big{)}$. Overall, this loss can fine-tune the adapter parameters $\Delta\boldsymbol{\phi}$ such that the SDS optimization pipeline in Eq. 1 can be used to smoothly manipulate our avatar by adjusting the scale factor $\alpha$ according to the user-input slider interpolation factor. Note that, with this loss design, we only need to identify the concept axis once (via LDA), and it can be used for the entire adapter fine-tuning process.
@@ 4.2 Preserving of Identity-related Attributes
While the Concept Sliding Loss presented above facilitates learning and representing intermediate points between opposing sides of a concept, it may inadvertently alter other identifying attributes of the avatar that we want to retain. This challenging issue arises as the 3D avatars contain many elements (e.g., face shape, hairstyle and clothing type) that involve entangled concepts such as gender, race, and age, as highlighted in challenge 2 in the introduction. Isolating and editing only the desired concept, while retaining the other identifying information of the avatar is challenging. To address this, we leverage Principal Component Analysis (PCA) to disentangle the concepts of the avatar and retain key identifying attributes. PCA helps to identify a small set of highly informative bases that are orthogonal to our concept axis, representing key identifying attributes unrelated to the target concept. Then, we propose an Attribute Preserving Loss $\mathcal{L}_{preserving}$ based on these key attribute bases found via PCA, enabling fine-tuning of the diffusion model adapter to preserve identifying attributes. Below, we first introduce the bases identification with PCA, followed by the attribute preserving loss.
@@
Attribute Bases Identification with PCA. Firstly, we reuse the features $\mathbf{F}_{p,i,x,y}$ and $\mathbf{F}_{n,i,x,y}$ that were computed in Sec. 4.1, which represent the positive and negative sides of the target concept. Then, we run PCA analysis on these features to identify a small set of key attributes in the feature space that are not related to our target concept. This is achieved through finding the $K$ principal components (i.e., orthogonal bases $\{\mathbf{b}_{a,k}\}_{k=1}^{K}$) which are orthogonal to our concept axis. Specifically, one way of performing PCA is through iteratively searching for the next principal component, which explains as much of the remaining variance of the data as possible, while being orthogonal to the previous identified principal components. Here, we recurrently employ this PCA technique to compute the main component vectors orthogonal to our concept axis, formulated as:
$\mathbf{b}_{a,k}=\mathop{\arg\max}\limits_{\mathbf{w}}\frac{\mathbf{w}\mathbf{\hat{F}}_{k}\mathbf{\hat{F}}_{k}^{T}\mathbf{w}^{T}}{\mathbf{w}\mathbf{w}^{T}},\text{where }\mathbf{\hat{F}}_{k}=\mathbf{\hat{F}}-\sum_{j=1}^{j-1}\mathbf{\hat{F}}\mathbf{b}_{a,j-1}^{T}\mathbf{b}_{a,k-1},$ (4)
@@
where $\mathbf{\hat{F}}\in\mathbb{R}^{D\times(2\cdot H\cdot W\cdot N_{S})}$ is the combined merger of both positive and negative sets of features ($\{\mathbf{F}_{p,i,x,y}\}_{i\in[1,N_{s}],x\in[1,H],y\in[1,W]}$ and $\{\mathbf{F}_{n,i,x,y}\}_{i\in[1,N_{s}],x\in[1,H],y\in[1,W]}$), and $\mathbf{w}$ is constrained to be a unit vector. Crucially, by treating the first ‘principal component’ $\mathbf{b}_{a,0}$ as our concept axis $\mathbf{b}_{c}$ identified in Eq. 2 and iteratively solving Eq. 4 for $k=\{1,...,K\}$, we can obtain $K$ key attribute bases $\{\mathbf{b}_{a,k}\}_{k=1}^{K}$ that best explain the remaining variance in the features. Furthermore, these obtained bases are orthogonal to our concept axis $\mathbf{b}_{c}$, which means they are key attributes that are unrelated to the target concept. Note that, although the small set of $K$ attribute bases has only a low rank, they capture the most important identifying information in the avatar that should be retained.
@@
Attribute Preserving Loss. With the identified attribute bases above that represent important key identifying attributes which should be retained, we introduce our attribute preserving loss to fine-tune the adapter ($\Delta\boldsymbol{\phi}$) to disentangle its editing effects, avoiding the editing of identifying attributes of the avatar. In other words, when users adjust the scale factor ($\alpha$) to perform editing, we expect the other key attributes to remain unchanged, as if the adapter is not applied for those other attributes. Thus, the attribute preserving loss $\mathcal{L}_{preserving}$ can be defined as:
@@
$\mathcal{L}_{preserving}(\Delta\boldsymbol{\phi})=\mathbb{E}_{\alpha\in[-1,1]}\Big{[}\sum_{k=1}^{K}\big{|}\big{|}\frac{\mathbf{f}(\alpha,\Delta\boldsymbol{\phi},c_{nu})\cdot\mathbf{b}_{a,k}}{\mathbf{b}_{a,k}\cdot\mathbf{b}_{a,k}}\mathbf{b}_{a,k}^{T}-\frac{\mathbf{f}(\varnothing,c_{nu})\cdot\mathbf{b}_{a,k}}{\mathbf{b}_{a,k}\cdot\mathbf{b}_{a,k}}\mathbf{b}_{a,k}^{T}\big{|}\big{|}\Big{]},$ (5)
@@
where $\mathbf{f}(\varnothing,c_{nu})$ represents the original latent features without applying the adapter. Intuitively, this loss compares the features after adaptation with the features that have not undergone adaptation, comparing them along a few key orthogonal dimensions that represent the key attributes of the avatar which should be retained. Importantly, this loss helps to optimize the adapter parameters $(\Delta\boldsymbol{\phi})$ to disentangle the effects of editing along the concept axis from the other main attributes and dimensions that are not related to the concept axis, encouraging the other attributes to be retained.
@@
Overall, by applying both the sliding loss $\mathcal{L}_{sliding}$ (Eq. 3) and the attribute preserving loss $\mathcal{L}_{preserving}$ (Eq. 5) to fine-tune the adapter $\Delta\boldsymbol{\phi}$ (see Fig. 2(a)), the model gains the ability to perform precise concept-specifc editing via the SDS optimization pipeline in Eq. 1. Thus, when the user manipulates the scale factor $\alpha$ at deployment time, the model can edit the 3D avatar towards a precise point along the concept axis, while maintaining the key identifying features.
4.3 Concept Sensitivity-Aware 3DGS Primitive Selection
The previous subsections introduced two loss criteria to fine-tune a concept-specific diffusion adapter, enabling precise control of avatar concepts through the SDS optimization pipeline. However, as mentioned in challenge 3, applying this slider-like control in practical applications often requires fast feedback from the model, to enable users to iteratively edit avatars towards their desired results, which necessitates high rendering and optimization speeds.
@@ 4.3 Concept Sensitivity-Aware 3DGS Primitive Selection
Therefore, to improve the rendering and optimization speed of our editing pipeline, we explore the 3DGS representation for fast 3D avatar editing. 3DGS is known for its efficiency in rendering and optimization [26, 21, 32], and has recently been of great interest for generation of 3D objects [37, 44, 7] and 3D humans [26]. Thus, considering its efficiency, we explore using 3DGS representation for our 3D avatar editing process. Specifically, the 3DGS representation consists of a set of $M$ Gaussian primitives which we denote as $\{\boldsymbol{\theta}_{i}\}_{i=1}^{M}$, where each $i$-th Gaussian primitive contains a set of parameters $\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\sigma_{i},\mathbf{h}_{i}\}$ that determine their location ($\boldsymbol{\mu}_{i}$), shape ($\boldsymbol{\Sigma}_{i}$), opacity ($\sigma_{i}$), and color ($\mathbf{h}_{i}$). During rendering, the color of each pixel is computed based on all the parameters of all $M$ primitives $\{\boldsymbol{\theta}_{i}\}_{i=1}^{M}$ in a differentiable manner, which facilitates the backpropagation of gradients.
@@
However, we find that optimizing the entire 3DGS representation (i.e., all $M$ primitives) yields sub-optimal results for 3D editing. This is because, not all primitives contribute to the final rendering equally [33], some of them are not so important, and so optimizing all primitives incurs high redundancy and harms optimization and editing efficiency, In addition, our concept-specific 3D editing task typically requires to modify only a small subset of primitives that contribute the most towards the target concept. Thus, we select and edit only a subset of primitives, as discussed below.
@@
Concept-Sensitive Primitive Selection. We propose a primitive selection mechanism based on concept-sensitivity to further improve the efficiency of the 3DGS representation in our editing pipeline. Instead of editing all $M$ 3DGS primitives, we only edit the primitives that contribute the most towards our target concept. To achieve this, our mechanism leverages the concept axis identified in Sec. 4.1 to quantify the contribution for each primitive, by computing a Concept Sensitivity score that measures how much each primitive is related to the target concept, i.e., how the changes in the primitive parameters affect the precise location along the concept axis. Then, by selecting primitives with a high Concept Sensitivity score and abediting only these primitives, we can effectively reduce the redundancy during the editing stage. We describe the details below.
@@
Firstly, to compute the sensitivity of Gaussian primitives with respect to the target concept (i.e., the identified concept axis $\mathbf{b}_{c}$ in the diffusion model’s latent space as discussed in Sec. 4.1), we need to express our 3D avatar in the same latent space. To do this, we first render our 3D avatar – which is represented by a set of primitives $\{\boldsymbol{\theta}_{i}\}_{i=1}^{M}$ – to $V$ randomly sampled views, yielding a set of rendered images. These rendered images are then passed into the diffusion model, to extract a set of latent features as $\{\mathbf{z}_{v,x,y}\}_{v\in[1,V],x\in[1,H],y\in[1,W]}$, where $\mathbf{z}_{v,x,y}\in\mathbb{R}^{D}$. These latent features $\mathbf{z}_{v,x,y}$ belong to the same latent space as $\mathbf{F}_{p,i,x,y}$ and $\mathbf{F}_{n,i,x,y}$ in Sec. 4.1.
@@
Next, we calculate the Concept Sensitivity score based on these latent features. Intuitively the Concept Sensitivity score measures the sensitivity of each primitive towards the target concept, i.e., how much movement there is along the identified concept axis $\mathbf{b}_{c}$ given a small change in the primitive’s parameters. To compute this, we first compute the degree of alignment $\bar{C}$ between latent features $\mathbf{z}_{v,x,y}$ and concept basis $\mathbf{b}_{c}$ via the dot product as follows: $\bar{C}=\sum_{v=1}^{V}\sum_{x=1}^{H}\sum_{y=1}^{W}\mathbf{b}_{c}\mathbf{z}_{v,x,y}^{T}$. The higher the value of $\bar{C}$, the more aligned the 3D human avatar is towards the target concept. Hence, if we take a gradient of $\bar{C}$ with respect to the primitive parameters, it will measure how much movement there is along the concept axis, given a small change in the primitive, i.e., the sensitivity of the target concept with respect to each primitive’s parameters. Specifically, for each $i$-th Gaussian primitive $\boldsymbol{\theta}_{i}$ with parameters $\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\boldsymbol{\theta}_{i},\mathbf{h}_{i}\}$, we compute its Concept Sensitivity score $S_{i}$ as:
$S_{i}=\sum_{p_{i}\in\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\sigma_{i},\mathbf{h}_{i}\}}\bigg{|}\frac{\partial\bar{C}}{\partial p_{i}}\bigg{|},$ (6)
@@
which sums the magnitude of the gradients w.r.t. each parameter. The higher the Concept Sensitivity score $S_{i}$, the more movement there will be along the concept axis given a small change in the $i$-th Gaussian primitive, which means that primitive is more crucial to editing of the concept. To select the most important primitives to edit, we pre-define a small fractional threshold $\gamma$, such that only the top $\gamma M$ primitives, i.e., $\{\boldsymbol{\theta}_{i}\}_{i=1}^{\gamma M}$, are selected for SDS optimization, instead of all $M$ primitives. Overall, this leads to much faster editing times, facilitating the fast feedback to users when using our ACS.
4.4 Implementation Details

### v3

@@ 4.1 Precise Slider-like Concept Editing
$\mathcal{L}_{CS}(\Delta\boldsymbol{\phi})=\mathbb{E}_{\alpha_{1},\alpha_{2}}\Big{[}\big{|}\big{|}(\operatorname{Proj}(\mathbf{f}(\alpha_{1},\Delta\boldsymbol{\phi},c_{nu}),\mathbf{b}_{c})-\operatorname{Proj}(\mathbf{f}(\alpha_{2},\Delta\boldsymbol{\phi},c_{nu}),\mathbf{b}_{c}))$
$-(\operatorname{Proj}\big{(}\frac{1+\alpha_{1}}{2}\boldsymbol{\mu}_{p}+\frac{1-\alpha_{1}}{2}\boldsymbol{\mu}_{n},\mathbf{b}_{c}\big{)}-\operatorname{Proj}\big{(}\frac{1+\alpha_{2}}{2}\boldsymbol{\mu}_{p}+\frac{1-\alpha_{2}}{2}\boldsymbol{\mu}_{n},\mathbf{b}_{c}\big{)})\big{|}\big{|}^{2}\Big{]},$ (3)
where $\mathbf{f}(\cdot)$ represents the generation of latent features, $c_{nu}$ is a neutral text prompt template, $\boldsymbol{\mu}_{p},\boldsymbol{\mu}_{n}\in\mathbb{R}^{D}$ are computed mean vectors for the positive and negative concept sides (see Supp for more details), and we select various points $\alpha_{1},\alpha_{2}$ in the range $[-1,1]$ to represent intermediate concepts; $\operatorname{Proj}(\cdot,\mathbf{b}_{c})$ is the projection operator, which projects the features onto the concept axes $\mathbf{b}_{c}$, where $\operatorname{Proj}(x,\mathbf{b}_{c})=\frac{x\cdot\mathbf{b}_{c}}{\mathbf{b}_{c}\cdot\mathbf{b}_{c}}\mathbf{b}_{c}^{T}$. Intuitively, this loss encourages the latent feature’s movements along the concept axes (after applying the slider scale factor) to match the reference movements as calculated with the aid of the positive and negative concept text prompts. The latter is computed using the interpolations between $\boldsymbol{\mu}_{n}$ and $\boldsymbol{\mu}_{p}$ as: $\big{(}\frac{1+\alpha_{1}}{2}\boldsymbol{\mu}_{p}+\frac{1-\alpha_{1}}{2}\boldsymbol{\mu}_{n}\big{)}$ and $\big{(}\frac{1+\alpha_{2}}{2}\boldsymbol{\mu}_{p}+\frac{1-\alpha_{2}}{2}\boldsymbol{\mu}_{n}\big{)}$. Moreover, our loss is applied to encourage precise movements of the slider, and do not penalize the actual locations of latent features on the concept axes. This minimizes “calibration” issues during training where the source avatar is not neutral, e.g., when an “old man” is given the source avatar, we do not assume that the source avatar is at scale factor of “0”.
@@
Overall, this loss can fine-tune the adapter parameters $\Delta\boldsymbol{\phi}$ such that the SDS optimization pipeline in Eq. 1 can be used to smoothly edit our avatar by adjusting the scale factor $\alpha$ during editing. Note that this loss can also be expanded to include more than two classes, where the dimensions of $\mathbf{b}_{c}$ are expanded, and concept means are also adapted to include the new features (e.g., neutral features). Also note that, with this loss design, we only need to identify the concept axes once (via LDA), and it can be used for the entire adapter fine-tuning process. Please refer to Supp for more details.
@@ 4.2 Preserving of Identity-related Attributes
While the Concept Sliding Loss presented above facilitates learning and representing intermediate points between opposing sides of a concept, it may inadvertently alter other identifying attributes of the avatar that we want to retain. This challenge arises as the 3D avatars contain many elements (e.g., face shape, hairstyle and clothing type) that involve entangled concepts such as gender, race, and age, as highlighted in challenge 2 in the introduction. Thus, it can be challenging to retain these other identifying information of the avatar while editing the desired concept.
To address this, we leverage Principal Component Analysis (PCA) to disentangle the concepts of the avatar and retain key identifying attributes. PCA helps to identify a small set of highly informative bases that are orthogonal to our concept axes, representing key identifying attributes unrelated to the target concept, but which can explain the other key variations in the latents. Based on these key attribute bases found via PCA, we propose an Attribute Preserving Loss $\mathcal{L}_{AP}$, enabling fine-tuning of the adapter to preserve identifying attributes. Below, we first discuss the bases identification with PCA, followed by the attribute preserving loss.
@@
Attribute Bases Identification with PCA. Firstly, we reuse the features ($\mathbf{F}_{p,i,x}$ and $\mathbf{F}_{n,i,x}$) that were computed in Sec. 4.1, which represent the positive and negative sides of the target concept. Then, we run PCA analysis on these features to identify a small set of key attributes in the feature space that are not related to our target concept. This is achieved through finding the $K$ principal components (i.e., orthogonal bases $\{\mathbf{b}_{a,k}\}_{k=1}^{K}$) which are orthogonal to our concept axes. Specifically, one way of performing PCA is through iteratively searching for the next principal component, which explains as much of the remaining variance of the data as possible, while being orthogonal to the previous identified principal components. Here, we recurrently employ this PCA technique to compute the main component bases orthogonal to our concept axis $\mathbf{b}_{c}$, formulated as:
$\mathbf{b}_{a,k}=\mathop{\arg\max}\limits_{\mathbf{w}}\frac{\mathbf{w}\mathbf{\hat{F}}_{k}\mathbf{\hat{F}}_{k}^{T}\mathbf{w}^{T}}{\mathbf{w}\mathbf{w}^{T}},\text{where }\mathbf{\hat{F}}_{k}=\mathbf{\hat{F}}-\sum_{j=1}^{k-1}\mathbf{\hat{F}}\mathbf{b}_{a,j-1}^{T}\mathbf{b}_{a,j-1},$ (4)
@@
where $\mathbf{\hat{F}}\in\mathbb{R}^{D\times(2\cdot N_{S}\cdot C)}$ is the combined merger of both positive and negative sets of features ($\{\mathbf{F}_{p,i,x}\}_{i\in[1,N_{s}],x\in[1,C]}$ and $\{\mathbf{F}_{n,i,x}\}_{i\in[1,N_{s}],x\in[1,C]}$), and $\mathbf{w}$ is constrained to be a unit vector. Crucially, by treating the first ‘principal component’ $\mathbf{b}_{a,0}$ as our concept axes $\mathbf{b}_{c}$ identified in Eq. 2 and iteratively solving Eq. 4 for $k=\{1,...,K\}$, we can obtain $K$ key attribute bases $\{\mathbf{b}_{a,k}\}_{k=1}^{K}$ that best explain the remaining variance in the features. Furthermore, these obtained bases are orthogonal to our concept axis $\mathbf{b}_{c}$ which are the key discriminative concept axes, which suggests that these bases are unrelated to the target concept. Note that, although the small set of $K$ attribute bases has only a low rank, they capture the most important identifying information in the avatar (that are unrelated to the target concept), which should be retained.
@@
Attribute Preserving Loss. With the identified attribute bases above that represent important key identifying attributes which should be retained, we introduce our attribute preserving loss to fine-tune the adapter ($\Delta\boldsymbol{\phi}$) to disentangle its editing effects, avoiding the editing of identifying attributes of the avatar. In other words, when users adjust the scale factor ($\alpha$) to perform editing, we expect the other key attributes to remain unchanged, as if the adapter is not applied for those other attributes. Thus, the attribute preserving loss $\mathcal{L}_{AP}$ can be defined as:
@@
$\mathcal{L}_{AP}(\Delta\boldsymbol{\phi})=\mathbb{E}_{\alpha}\Big{[}\sum_{k=1}^{K}\big{|}\big{|}\operatorname{Proj}(\mathbf{f}(\alpha,\Delta\boldsymbol{\phi},c_{nu}),\mathbf{b}_{a,k})-\operatorname{Proj}(\mathbf{f}(\varnothing,c_{nu}),\mathbf{b}_{a,k})\big{|}\big{|}^{2}\Big{]},$ (5)
@@
where $\mathbf{f}(\varnothing,c_{nu})$ represents the original latent features without applying the adapter, and various values of $\alpha$ are selected in the range $[-1,1]$. Intuitively, this loss compares the features after slider adaptation with the features that have not undergone adaptation, comparing them along a few key orthogonal dimensions that represent the key attributes of the avatar which should be retained. Importantly, this loss helps to optimize the adapter parameters $(\Delta\boldsymbol{\phi})$ to disentangle the effects of editing along the concept axis from the other main attributes and dimensions that are not related to the concept axis, encouraging the other attributes to be retained.
@@
Overall, by applying both the sliding loss $\mathcal{L}_{CS}$ (Eq. 4.1) and the attribute preserving loss $\mathcal{L}_{AP}$ (Eq. 5) to fine-tune the adapter $\Delta\boldsymbol{\phi}$ (see Fig. 2(a)), the model gains the ability to perform precise concept-specifc editing via the SDS optimization pipeline in Eq. 1. Thus, when the user manipulates the scale factor $\alpha$ at deployment time, the model can edit the 3D avatar towards a precise point along the concept axis, while maintaining the key identifying features.
4.3 Concept-sensitive 3DGS Primitive Selection
The previous subsections introduced two loss criteria to fine-tune a concept-specific diffusion adapter, enabling precise control of avatar concepts through the SDS optimization pipeline. Here, we further explore a technique to reduce redundant computations, further improving the efficiency of 3D human avatar editing, specifically for avatars in the 3D Gaussian Splatting (3DGS) representation.
@@ 4.3 Concept-sensitive 3DGS Primitive Selection
3DGS is known for its efficiency in rendering and optimization [21, 17, 26], and has recently been of great interest for generation of 3D objects [33, 38, 6] and 3D humans [21]. Thus, considering its efficiency, we adopt the 3DGS representation for our 3D avatars. The 3DGS representation consists of a set of $M$ Gaussian primitives which we denote as $\{\boldsymbol{\theta}_{i}\}_{i=1}^{M}$, where each $i$-th Gaussian primitive contains a set of parameters $\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\sigma_{i},\mathbf{h}_{i}\}$ that determine their location ($\boldsymbol{\mu}_{i}$), shape ($\boldsymbol{\Sigma}_{i}$), opacity ($\sigma_{i}$), and color ($\mathbf{h}_{i}$). During rendering, the color of each pixel is computed based on all the parameters of all $M$ primitives $\{\boldsymbol{\theta}_{i}\}_{i=1}^{M}$ in a differentiable manner, which facilitates the backpropagation of gradients.
@@
However, we find that optimizing the entire 3DGS representation (i.e., all $M$ primitives) yields sub-optimal results for 3D editing. This is because, not all primitives contribute to the final rendering equally [27], some of them are not so important, and so optimizing all primitives incurs high redundancy and harms optimization and editing efficiency, In addition, our concept-specific 3D editing task typically needs to modify only a small subset of primitives that contribute the most towards the target concept. Thus, we select and edit only a subset of primitives, as discussed below.
@@
Concept-Sensitive Primitive Selection. We propose a primitive selection mechanism based on concept-sensitivity to further improve the efficiency of the 3DGS representation in our editing pipeline. Instead of editing all $M$ 3DGS primitives, we only edit the primitives that contribute the most towards our target concept. To achieve this, our mechanism leverages the concept axes identified in Sec. 4.1 to quantify the contribution for each primitive, by computing a Concept Sensitivity score that measures how much each primitive is related to the target concept, i.e., how the changes in the primitive parameters affect the precise location along the concept axis. Then, by selecting primitives with a high Concept Sensitivity score and edit only these primitives, we can effectively reduce the redundancy during the editing stage.
@@
Firstly, to compute the sensitivity of Gaussian primitives with respect to the target concept (i.e., the identified concept axis $\mathbf{b}_{c}$ in the diffusion model’s latent space as discussed in Sec. 4.1), we need to express our 3D avatar in the same latent space. To do this, we first render our 3D avatar – which is represented by a set of primitives $\{\boldsymbol{\theta}_{i}\}_{i=1}^{M}$ – to $V$ randomly sampled views, yielding a set of rendered images. These rendered images are then passed into the diffusion model, to extract a set of latent features as $\{\mathbf{z}_{v,x,y}\}_{v\in[1,V],x\in[1,C]}$, where $\mathbf{z}_{v,x}\in\mathbb{R}^{D}$. These latent features $\mathbf{z}_{v,x}$ belong to the same latent space as $\mathbf{F}_{p,i,x}$ and $\mathbf{F}_{n,i,x}$ in Sec. 4.1.
@@
Next, we calculate the Concept Sensitivity score based on these latent features. Intuitively, the Concept Sensitivity score measures the sensitivity of each primitive towards the target concept, i.e., how much movement there is along the identified concept axis $\mathbf{b}_{c}$ given a small change in the primitive’s parameters. To compute this, we first compute the projection of the features $\mathbf{z}_{v,x}$ onto the concept bases $\mathbf{b}_{c}$ as follows: $\bar{C}=\frac{1}{V}\sum_{v=1}^{V}\operatorname{Proj}(\mathbf{z}_{v},\mathbf{b}_{c})$. The higher the values in $\bar{C}$, the more aligned the 3D human avatar is towards the target concept. Hence, if we take a gradient of $\bar{C}$ with respect to the primitive parameters, it will measure how much movement there is along the concept axes, given a small change in the primitive, i.e., the sensitivity of the target concept with respect to each primitive’s parameters. Specifically, for each $i$-th Gaussian primitive $\boldsymbol{\theta}_{i}$ with parameters $\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\boldsymbol{\theta}_{i},\mathbf{h}_{i}\}$, we compute its Concept Sensitivity score $S_{i}$ as:
$S_{i}=\sum_{p_{i}\in\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i},\sigma_{i},\mathbf{h}_{i}\}}\sum_{l=1}^{L}\bigg{|}\frac{\partial\bar{C}_{l}}{\partial p_{i}}\bigg{|},$ (6)
@@
which sums the magnitude of the gradients w.r.t. each parameter, where $L$ is the dimensionality of $\bar{C}$. The higher the Concept Sensitivity score $S_{i}$, the more movement there will be along the concept axes given a small change in the $i$-th Gaussian primitive, which means that primitive is more crucial to editing of the concept. To select the most important primitives to edit, we pre-define a small fractional threshold $\gamma$, such that only the top $\gamma M$ primitives, i.e., $\{\boldsymbol{\theta}_{i}\}_{i=1}^{\gamma M}$, are selected for SDS optimization, instead of all $M$ primitives. Overall, this leads to significantly improved efficiency.
5 Implementation Details